### v1

# Leveraging Large Language Models for Explainable Activity Recognition in Smart Homes: A Critical Evaluation

Abstract.
Explainable Artificial Intelligence (XAI) aims to uncover the inner reasoning of machine learning models. In IoT systems, XAI improves the transparency of models processing sensor data from multiple heterogeneous devices, ensuring end-users understand and trust their outputs. Among the many applications, XAI has also been applied to sensor-based Activities of Daily Living (ADLs) recognition in smart homes. Existing approaches highlight which sensor events are most important for each predicted activity, using simple rules to convert these events into natural language explanations for non-expert users. However, these methods produce rigid explanations lacking natural language flexibility and are not scalable. With the recent rise of Large Language Models (LLMs), it is worth exploring whether they can enhance explanation generation, considering their proven knowledge of human activities. This paper investigates potential approaches to combine XAI and LLMs for sensor-based ADL recognition. We evaluate if LLMs can be used: a) as explainable zero-shot ADL recognition models, avoiding costly labeled data collection, and b) to automate the generation of explanations for existing data-driven XAI approaches when training data is available and the goal is higher recognition rates. Our critical evaluation provides insights into the benefits and challenges of using LLMs for explainable ADL recognition.
Smart Homes, Human Activity Recognition, XAI, LLMs
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TIOTccs: Human-centered computing Ambient intelligenceccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Natural language generation
This manuscript is under review. Please refer to authors for up-to-date information.
1. Introduction
The Internet of Things (IoT) and pervasive computing communities have investigated sensor-based Human Activity Recognition (HAR) for decades, designing and testing new recognition methods by exploiting the evolution of sensing devices (Babangida et al., 2022). Among the many applications of HAR, the recognition of Activities of Daily Living (ADL) in smart homes may have high-impact healthcare applications, including the early detection and continuous monitoring of cognitive decline (Bakar et al., 2016).
Most existing smart home ADL recognition approaches are based on data-driven deep learning models that require large datasets of (possibly labeled) sensor events. The lack of suitable datasets, especially for the smart home environments, is known as the data scarcity problem (Tang et al., 2021), and together with subject heterogeneity (Soleimani and Nazerfard, 2021), and environment and sensing infrastructure heterogeneity (Chiang et al., 2017) is one of the reasons of the limited generalization ability of current models. An additional and not secondary problem is the opacity of deep learning models in their activity prediction process. A few works in recent years proposed the adoption of eXplainable Artificial Intelligence (XAI) techniques to provide the rationale of deep learning ADLs recognition models (Arrotta et al., 2022a; Das et al., 2023). Several decisions in pervasive computing applications may rely on activity prediction, and inferring why an ADL was predicted may lead to more understandable, trusted, and transparent systems (Wolf, 2019). This is the case for applications that support clinicians in their diagnosis process (Lussier et al., 2018). Existing XAI methods for sensor-based ADL recognition adopt heuristic algorithms to generate sentences in natural language by considering the sensor events in the input time window that were more relevant for the prediction (Das et al., 2023; Arrotta et al., 2022a) but the generalization capabilities of those heuristics is unclear. Indeed, these approaches generate rigid explanations that lack the nuance and flexibility of natural language and, at the same time, incorporating new activities or devices would require the significant effort of manually adding new rules.
In this paper, we critically evaluate the role that LLMs may potentially have in explainable sensor-based HAR (XAR). There have been preliminary investigations of applying LLMs to HAR (Cleland et al., 2024), leveraging the common-sense knowledge about human activities encoded in LLMs to reduce the required labeled data to train supervised classifiers. However, to the best of our knowledge, no existing work studies how LLMs can also be leveraged for explanations and the possible dangers they may introduce. We aim to answer these research questions:
Q1) Is it possible to develop a zero-shot ADL recognition method (i.e., not requiring any training data) based on LLM that also provides meaningful explanations?
Q2) Is it possible to use LLM to automatically generate explanations in natural language, independently from the underlying XAI method adopted?
Q3) What are the drawbacks and risks in using LLM for XAR?
Our results answer positively to Q1 by implementing an end-to-end LLM-based system and testing it on two public datasets, comparing recognition accuracy with a deep learning baseline and an LLM method that does not provide explanations. We also positively answer to Q2 by proposing a standard format that different XAI methods may adopt to represent the most relevant features leading to the prediction. By using this input and an appropriate prompting technique, LLMs can generate appreciated explanations as proved by user surveys. Regarding Q3, we discuss the potential disadvantages of such systems and consider possible mitigation actions. We identify over-reliance as the main risk, by providing examples of intuitive LLM explanations for wrong predictions. We also discuss cost and privacy issues, real-life deployment concerns, and the possible impact of LLMs’ hallucinations.
In summary, the main contributions are the following:
  • We propose LLMe2e: a zero-shot LLM method for ADL recognition that also provides human-readable explanations.
  • We propose LLMExplainer: an LLM-based approach to generate explanations in natural language for ADL recognition in smart homes. LLMExplainer is based on a standardized format to represent the output of XAR methods and, therefore, is agnostic to the specific XAR model being adopted.
  • We conduct extensive experiments showing that LLMe2e is slightly less accurate than a state-of-the-art baseline while still providing adequate recognition rates. Based on two user surveys with more than $200$ subjects, we show that LLM-based approaches offer explanations that users appreciate more than the ones generated by state-of-the-art heuristic-based methods.
  • While the results are promising, our critical evaluation highlights the risks and the drawbacks of adopting LLMs in XAR, including examples illustrating the risk of over-reliance on explanations.
2. Related Work
2.1. Explainable Human Activity Recognition
Several eXplainable Artificial Intelligence (XAI) approaches have been proposed to explain the output of smart home ADL recognition based on deep learning models (Arrotta et al., 2022a, b; Das et al., 2023). For instance, the work in (Das et al., 2023) uses post hoc XAI methods such as LIME (Ribeiro et al., 2016) and SHAP (Scott et al., 2017) to derive the most important sensor events that an LSTM model used to predict the activity. Post hoc methods consider the deep learning model as a black box, and they analyze the correlations between the classifier’s inputs and outputs to determine the most important features for the prediction. On the other hand, the work in (Arrotta et al., 2022a) adopts a convolutional neural network designed to be explainable (Rudin, 2019), learning prototypes of input data during training and using them to explain the decision process during classification. In these works, explanations in natural language are generated with heuristic approaches (i.e., simple rules) and describe which sensor events the classifier considers important to perform a prediction. Such approaches are quite rigid in generating explanations, lacking the nuance and flexibility of natural language (especially with long explanations). At the same time, they lack scalability: incorporating new activities or devices would require manually adding new rules. To the best of our knowledge, LLMs have never been explored as a flexible method to generate explanations for ADLs recognition. A major challenge in XAI is also to quantitatively evaluate the effectiveness of the generated explanations (Mohseni et al., 2021). The most common method is to perform user surveys (Das et al., 2023; Arrotta et al., 2022a; Jeyakumar et al., 2023) and we follow this approach. An interesting recent work shows that LLMs could also be used to evaluate explanations, and can provide results very close to those of a survey (Fiori et al., 2024).
2.2. Large Language Models for Human Activity Recognition
The HAR community is actively working on leveraging Large Language Models (LLMs). Several existing approaches take advantage of the intrinsic knowledge of LLMs about human activities to reduce the amount of required training data (Liu et al., 2023; Okita et al., 2023; Zhou et al., 2023; Zilelioglu et al., 2023; Arrotta et al., 2024; Thukral et al., 2024; Leng et al., 2024). Most interestingly, some recent works (most of them are still pre-prints) considered the more challenging setting of using LLMs for zero-shot recognition of human activities (Ji et al., 2024; Hota et al., 2024). These works leverage a textual representation of input sensor data to query the LLM for the activity most likely performed by the monitored subject. Several research papers recently proposed LLM-based zero-shot approaches for ADL recognition in smart homes, proposing specifically designed prompts enabling LLMs to reason on textual representation of input sensor data to identify the most likely ADLs performed by the inhabitants (Gao et al., 2024; Cleland et al., 2024; Xia et al., 2023; Civitarese et al., 2024; Chen et al., 2024; Fritsch et al., 2024). However, all of these works focus only on recognizing the performed activities, without providing explanations that non-expert users may leverage to grasp the rationale behind the decision.
2.3. Combining Large Language Models and Explainable AI
There are only a few works that investigated the role of LLMs in improving eXplainable AI. Note that, in this work, we are not interested in the more challenging task of explaining LLMs output (Zhao et al., 2024), but in the use of LLMs to support explainable AI in the HAR tasks. For instance, the studies in (Ali and Kostakos, 2023; Mavrepis et al., 2024) indicate that LLMs could take the output of XAI approaches (i.e., the prediction and the feature importance) and generating natural language explanations. However, to the best of our knowledge, this approach was never studied in the HAR domain.
3. LLM-based methods for Explainable ADLs Recognition
In this section, we present two novel approaches exploring how LLMs can be adopted for XAI in sensor-based ADL recognition. In Section 5, we will also provide a critical evaluation of these methods in subsequent sections to assess their strengths, limitations, and potential implications.
3.1. The Explainable Human Activity Recognition Problem
We first formally define the XAR problem that we aim to address in this work. Like many studies in this field, we focus on a smart home setting where either a single individual resides or the sensing system is capable of accurately linking each sensor activation to the specific subject responsible for triggering it. The home is equipped with several environmental sensors that capture the interaction of the subject with the surrounding environment and the person may be carrying a wearable device revealing low-level physical activities. Depending on its type, each sensor generates a data stream corresponding to activation/deactivation (e.g., open/close), measurements (e.g., humidity values), or inertial sensor data. Measurements and inertial data are preprocessed to generate higher-level information (e.g., high humidity or sitting posture) so that each sensor has an associated time series of events that identify the start and end time of the detected conditions (e.g., high humidity start/end, sitting start/end). The multi-variate time series corresponding to the whole set of sensors is segmented in temporal windows.
Given a window $\mathbf{w}$ in the time series of sensor events, and a set of candidate activities $\mathcal{A}=\{A_{1},A_{2},\dots,A_{n}\}$, the goal of HAR is to predict the most likely activity $\hat{A}\in\mathcal{A}$ that the subject is performing during the time of that window. This task is usually performed by a classifier $h$ such that $h\left(\mathbf{w}\right)=\hat{A}$. Explainable classifiers generally associate a value to each event in $\mathbf{w}$, denoting its importance for the prediction $h\left(\mathbf{w}\right)$. For each prediction $\hat{A}=h\left(\mathbf{w}\right)$, an explanation $\mathbf{e}(h\left(\mathbf{w}\right))$ is defined as the set of events in $\mathbf{w}$ considered by $h$ as the most important for its prediction (e.g., based on a threshold on the importance values).
For the sake of interpretability, the sequence of events $\mathbf{e}(h(\mathbf{w}))$ is usually reported in various human-readable forms including graphics or natural language, making explicit the type of events, their temporal relationships.
3.2. LLMe2e: Zero-Shot Explainable ADL Recognition
Starting from the work initially proposed in (Civitarese et al., 2024), we designed a novel Zero-Shot Explainable ADL recognition method based on LLMs. We will refer to this approach as LLMe2e. This method adopts an end-to-end approach, starting with sensor data and using a single LLM prompt to perform both ADL classification and natural language explanation generation simultaneously. LLMe2e does not require any training data, since it only leverages the intrinsic knowledge about human activities encoded in LLMs. The main differences with the work in (Civitarese et al., 2024) are a different representation of the input sensor data and the generation of explanations for each prediction. Note that even other zero-shot approaches proposed in the literature (e.g.,  (Cleland et al., 2024)) could be similarly extended to generate explanations. The high-level architecture of LLMe2e is shown in Figure 1.
Figure 1. LLMe2e: Zero-shot Explainable Activities of Daily Living Recognition
First, the raw stream of sensor data is transformed into a higher-level representation, referred to as sensor states, by the Sensor States Generation module. Intuitively, each sensor state represents the time interval during which a specific property detected by a sensor (e.g., the front door being opened) is true.
Next, the stream of sensor states is divided into fixed-time windows by the Segmentation module. Each window is then converted into a structured text format by the States2JSON module. Figure 2 shows a JSON standard format that we designed for sensor state representation. This format is used to provide the input to LLMe2e for activity prediction, but is also adopted by the LLMExplainer approach to represent the states corresponding to important events identified by XAR techniques, as we will discuss in Section 3.3.
Figure 2. JSON standard representation of the sensor states.
This textual data is used by the User Prompt Generation module to create the input for the large language model (LLM). The LLM is then queried with a system prompt (i.e., a detailed description of the ADL recognition and explanation generation tasks), along with the user prompt. Finally, the LLM’s output is parsed by the Activity Label and Explanation Extractor, which extracts the most likely performed ADL and provides the corresponding explanation.
3.2.1. Sensor State Generation
[…]
3.2.2. Segmentation
The Segmentation module segments the stream of sensor states into fixed-length time windows of $\tau$ seconds, with an overlap factor of $o$ between consecutive windows. Each time window is associated with the sensor states that are active during that period, but only considering the overlapping portions of the original sensor state intervals.
3.2.3. System Prompt
A system prompt has the role of instructing the LLM on the specific task(s) that it has to perform. The LLMe2e’s system prompt adopts a “role prompting” strategy, instructing the LLM to operate as a human activity recognition system. This strategy showed to be particularly effective for zero-shot tasks (Kong et al., 2023). The system prompt includes all the general information that apply to all the windows, in particular: a) information about the layout of the smart home and about the interactions between the subject and the home as they can be captured by the sensing infrastructure; b) the JSON input format and how to interpret it; c) the ADL classification task; d) the explanation generation task.
Figure 3. LLMe2e: template of the system prompt.
3.2.4. User Prompt Generation
Each user prompt defines the specific instance of the problem that the LLM has to process. In LLMe2e, the User Prompt Generation generates a user prompt by leveraging the JSON representation of the input window created by the States2JSON module, also appending instructions to adopt the well-known Chain of Thought (CoT) prompting strategy (Wei et al., 2022). An example of a user prompt is shown in Figure 4.
Figure 4. LLMe2e: An example of user prompt.
Figure 5 depicts an example of output generated by LLMe2e based on the system and user prompts illustrated above.
Figure 5. LLMe2e: output example
3.3. LLMExplainer: LLMs as explanation generators for data-driven XAR
In this Section, we propose LLMExplainer: an LLM-based method to generate natural language explanations from the most important events $\mathbf{e}$ derived by a data-driven XAR methods. LLMExplainer is agnostic to the underlying XAR methods being adopted, since the only requirement is that their output, i.e., the most important events contributing to each prediction, can be represented in the JSON format depicted in Figure 2. Considering the existing XAR methods, the conversion in this format is straightforward. The architecture of LLMExplainer is depicted in Figure 6.
Figure 6. LLMExplainer: Overall architecture.
3.3.1. System Prompt
While the system prompt of LLMExplainer (depicted in Figure 7) resembles the one of LLMe2e, it is only focused on the explanation generation task. Specifically, it includes a) a high-level description of the task, using the role-prompting strategy; b) the JSON input format representing the most important states identified by the explainable data-driven classifier and how to interpret it; c) instructions on how to generate the explanation; d) instructions to make sure that the explanations are suitable to non-expert users. Note in particular the sentence ”you can reason on the correlation between the states and on the temporal correlations among them, but do not add inferences that are not directly supported by data”. We added this part to better guide the model to stick to the facts stated in the user message.
Figure 7. LLMExplainer: System prompt.
3.3.2. User Prompt
The User Prompt Generation module combines the predicted activity and the most important features (formatted using our JSON format) produced by the XAR classifier on a specific window. An example of a user prompt and the corresponding output are shown in Figures 8 and 9.
Figure 8. LLMExplainer: Example of an user prompt.
Figure 9. LLMExplainer: Output example
As mentioned above, LLMExplainer can be easily used with any XAR model. However, an additional module converting the most important features derived by a specific XAR model into the JSON format is required. Indeed, LLMExplainer expects important features expressed as states. Each state should be associated with a name (e.g., “the stove is on”), and the list of intervals in which that state was active. In order to clarify how LLMExplainer generates the user prompt, we show an example using DeXAR (Arrotta et al., 2022a), representing a state-of-the-art XAR method. The application of LLMExplainer to other XAR approaches like the one presented in (Das et al., 2023) is straightforward. Figure 10 shows an explanation (i.e., a heatmap) obtained by DeXAR’s classifier when predicting the activity preparing hot meal.
Figure 10. An explanation of the DeXAR’s (Arrotta et al., 2022a) classifier for the predicted activity preparing hot meal.
Intuitively, each row of the heatmap encodes a condition detected by a sensor, and each segment in a row encodes the time interval when the condition was true. In DeXAR, the most important features are the rows where at least one pixel’s intensity is above a threshold. In order to create the JSON representation, a dedicated module converts each important feature name into a state name and, for each segment, uses its start and end time to represent the time intervals where the state was observed. Figure 11 shows how the most important features from the heatmap in Figure 10 would be converted into the JSON format to be used as input for LLMExplainer.
Figure 11. The user prompt for LLMExplainer generated from the most important features in Figure 10.
4. Experimental Evaluation
In this section, we describe how we evaluated LLMe2e and LLMExplainer and report experimental results.
4.1. Datasets
For the sake of this work, we selected two publicly available smart-home datasets that include several types of environmental sensors, with one of them including also inertial sensors from a smartwatch. We excluded datasets based almost entirely on PIR sensors, like CASAS (Cook et al., 2009) for reasons that are discussed in Section 5.
4.1.1. UCI ADL
The UCI ADL (Ordónez et al., 2013) dataset has been collected in two different homes (Home A and Home B), each one inhabited by a single subject. The dataset includes $14$ days of ADLs for Home A and $21$ days for Home B. Both homes are equipped with several types of sensors: movement (PIR), magnetic, pressure, and plug sensors. Unlike Home A, Home B has PIR sensors also installed over some doors. In contrast, Home A has a PIR sensor above the stove, a plug sensor to detect toaster usage, and a magnetic sensor to monitor interactions with the bathroom cabinet, which are not present in Home B. In both homes data was collected for the following activities: preparing breakfast, preparing lunch, snacking, personal care (i.e., grooming in the dataset), showering, leaving home, relaxing on couch (i.e., spare time/TV in the dataset), sleeping, and toileting. Home B includes an additional activity, that is preparing dinner. In our experiments, we did not consider the activity toileting as the semantic meaning of the label and its relation to the flush sensor is unclear. Since the sleeping and relaxing on the couch are over-represented in the dataset (due to a long duration compared to the other ADLs) and exhibit only a few unique sensor patterns, we down-sample these classes in order to have a balanced dataset.
4.1.2. MARBLE
The MARBLE (Arrotta et al., 2021) dataset includes data from $12$ subjects in both single and multi-inhabitant scenarios. For the sake of this work, we only consider the single-inhabitant scenarios of the dataset. The data was collected using smart plugs (to monitor stove and TV usage), magnetic (to monitor the interaction with drawers and cabinets), and pressure sensors (to monitor the usage of dining room chairs and the couch in the living room). Additionally, each subject wears a smartwatch to capture inertial data generated by hand gestures. Consistently with the work in (Arrotta et al., 2022a), we only differentiate between static and dynamic gestures. To detect incoming and outgoing phone calls, an Android application was installed on the smartphone of each subject. The original dataset includes 13 activities, but, following (Civitarese et al., 2024), we consider the following $11$ activities: clearing table, eating, entering home, leaving home, phone call, preparing cold meal, preparing hot meal, setting up table, taking medicines, using pc, and watching tv.
4.2. Baselines
In this work, we want to evaluate two aspects: 1) the recognition rate of LLMe2e, and 2) the quality of explanations of both LLMe2e and LLMExplainer. We consider the following two baselines:
  • DeXAR (Arrotta et al., 2022a): a fully supervised approach transforming sensor data into semantic images to leverage Convolutional Neural Networks and XAI methods. Although other baselines could have been considered (Das et al., 2023), this particular one was the only option for which we had access to the code necessary to conduct our experiments. We use the DeXAR implementation leveraging the XAI Model Prototypes approach (Rudin, 2019), which was considered the best one in the paper. Note that we slightly modified the implementation by removing the extraction of past activities. This is because the approach proposed in (Arrotta et al., 2022a) can not be applied to LLMe2e, since it would require the confidence of the activity prediction that cannot be provided by the LLM. In our experiments, we use DeXAR both as a baseline for explanations and as the XAR method for which LLMExplainer generates explanations.
  • ADL-LLM (Civitarese et al., 2024): a recent zero-shot ADL recognition method based on LLMs. We consider this method as a baseline for the recognition rate, as our LLMe2e approach is an extension and adaptation of it. Differently from LLMe2e, each window is translated into a sentence in natural language with a heuristic approach, rather than providing just the information of sensors’ triggering events like in the JSON template used in LLMe2e. ADL-LLM does not explicitly generate explanations, but we consider it to compare the recognition rate.
4.3. Experimental Setup
We developed the prototypes of LLMe2e and LLMExplainer in Python, querying the OpenAI ‘gpt-4o’ LLM model through the official OpenAI library, setting the temperature parameter to 00.
4.3.1. Pre-processing and dataset splitting for ADL recognition
Considering the segmentation window, we used parameters suggested in the literature. For the MARBLE dataset, we used 16-second windows with 80% overlap (Arrotta et al., 2021), while for the UCI ADL dataset, we used 60-second windows with 80% overlap (Ordónez et al., 2013). Since our approach is zero-shot and does not require training, we could ideally report results using the entire dataset. However, since in DeXAR the dataset was divided into $70\%$ for training and $30\%$ for testing (Arrotta et al., 2022a), we evaluate LLMe2e using the same test set. The test set for ADL-LLM was $70\%$ and we compare LLMe2e on the same set.
4.3.2. User Surveys
To assess the quality of the explanations from the point of view of non-expert users, we performed two user-based surveys: one for the MARBLE dataset and another for the Home B of the UCI ADL dataset. We excluded Home A from this analysis since it is a significantly simpler dataset compared to Home B (as also reflected by our results in Section 4.4). Frequently, in this dataset there is a only predominant state in the time windows, leading to relatively simplistic explanations.
Each survey includes, for each activity class, an instance of a window classification associated with the explanations of DeXAR, LLMe2e, and LLMExplainer for that same prediction. Note that in our surveys we included only correct classifications. In our implementation, LLMExplainer generates explanations using the result of the XAR technique implemented by DeXAR, as described in Section 3.3. Survey participants rated each explanation using the Likert scale. An example of a question in our survey is shown in Figure 12.
Figure 12. A sample question from our survey
[…]
4.4. Results
4.4.1. Recognition rate
Tables 1 and  2 compare the recognition rates of LLMe2e (not requiring any training data) and DeXAR (using a training dataset) measured by the weighted F1-score.
Table 1. MARBLE: F1 score of LLMe2e vs DeXAR on the same test set (i.e., $30\%$ of the dataset)
Activity DeXAR LLMe2e
Clearing table 0.36 0.37
Eating 0.96 0.90
Entering home 0.92 0.69
Leaving home 0.90 0.80
Phone call 0.98 0.82
Preparing cold meal 0.59 0.54
Preparing hot meal 0.86 0.83
Setting up table 0.60 0.23
Taking medicines 0.43 0.29
Using pc 0.96 0.96
Watching TV 0.99 0.90
F1 Weighted avg. 0.86 0.80
Table 2. UCI ADL: F1 score of LLMe2e vs DeXAR on the same test set (i.e., $30\%$ of the dataset)
Home A Home B
Activity DeXAR LLMe2e DeXAR LLMe2e
Leaving home 1.00 1.00 0.87 0.87
Personal care 0.99 1.00 0.93 0.95
Preparing breakfast 0.93 0.96 0.32 0.61
Preparing dinner 0.16 0.18
Preparing lunch 0.95 0.94 0.30 0.19
Relaxing on couch 0.99 0.99 0.88 0.76
Showering 0.99 1.00 0.91 0.93
Sleeping 1.00 1.00 0.99 0.96
Snacking 0.00 0.40 0.31 0.51
F1 Weighted Avg. 0.96 0.97 0.72 0.75
Considering MARBLE, DeXAR achieves a weighted F1-score $6\%$ higher than LLMe2e indicating the superiority of a method that uses training data. However, LLMe2e still reaches an acceptable weighted F1-score of $0.8$. LLMe2e performs significantly worse on the leaving home, entering home, setting up table, and taking medicines activities. More details about the recognition errors can be found in the confusion matrix in Figure 13.
Figure 13. LLMe2e: Confusion Matrix on MARBLE dataset
In general, the LLM has lower rates for activities that can be learned in a data-driven way by looking at the sensor patterns, while can be easily confused with others by only considering common-sense knowledge. For instance, LLMe2e often misclassifies setting up the table as clearing the table, preparing a cold meal, preparing a hot meal, or eating due to the similarity between these activities, involving interactions with kitchen items and frequent movement between the kitchen and dining room.
Considering the UCI ADL dataset, we observe that the recognition rates of LLMe2e and DeXAR are similar. An interesting insight is that, in Home A, DeXAR can not capture the Snacking activity, since it is poorly represented in the training set. On the other hand, even if with a low recognition rate, LLMe2e can capture this activity because it does not rely on a training dataset. More details about misclassifications can be found in the confusion matrices in Figure 14.
(a) Home A
(b) Home B
Figure 14. LLMe2e: Confusion matrices on UCI ADL dataset
For both homes, the main source of misclassification is represented by activities related to eating, since they involve similar sensor patterns.
Tables 3 and 4 compare the recognition rate of ADL-LLM and LLMe2e measured by the weighted F1-score [^3] For the sake of fairness, we did not consider dynamic manipulations as input for LLMe2e since these were not considered in the ADL-LLM paper (Civitarese et al., 2024).. The results show that LLMe2e achieves a similar recognition rate for both datasets despite the lower-level input representation (JSON format) and the additional task of generating explanations.
Table 3. MARBLE: F1 score of LLMe2e vs ADL-LLM on the same test set (i.e., $70\%$ of the dataset)
Activity ADL-LLM LLMe2e
Clearing table 0.25 0.14
Eating 0.91 0.90
Entering home 0.41 0.67
Leaving home 0.71 0.81
Phone call 0.97 0.82
Preparing cold meal 0.51 0.52
Preparing hot meal 0.82 0.82
Setting up table 0.13 0.25
Taking medicines 0.47 0.31
Using pc 0.97 0.96
Watching TV 0.97 0.93
F1 Weighted avg. 0.80 0.80
Table 4. UCI ADL: F1 score of LLMe2e vs ADL-LLM on the same test set (i.e., $70\%$ of the dataset)
Home A Home B
Activity ADL-LLM LLMe2e ADL-LLM LLMe2e
Leaving home 1.00 1.00 0.90 0.78
Personal care 0.97 0.99 0.96 0.95
Preparing breakfast 0.75 0.90 0.71 0.75
Preparing dinner 0.33 0.24
Preparing lunch 0.91 0.94 0.40 0.43
Relaxing on couch 0.96 0.99 0.71 0.81
Showering 0.99 1.00 0.95 0.96
Sleeping 1.00 1.00 0.94 0.99
Snacking 0.36 0.32 0.28 0.30
F1 Weighted Avg. 0.94 0.96 0.75 0.77
4.4.2. Explanation evaluation
In the following, we show the results of the user surveys. The box plots in Figure 15 depict the distribution of the scores obtained by DeXAR, LLMe2e, and LLMExplainer applied to the output of the DeXAR classifier.
(a) MARBLE
(b) UCI ADL Home B
Figure 15. Distribution of the scores provided by the participants of the user survey
For both datasets, the participating subjects gave a good score to the explanations generated by the LLM methods with LLMExplainer being the best, while they gave lower scores to the explanations produced by DeXAR. Even though these two methods leverage the same relevant sensor data (i.e., LLMExplainer generates explanations starting from the important events identified by DeXAR’s classifier), the LLM uses a more convincing wording including possible relationships between the states detected by sensors and the specific actions done by the inhabitant to perform the predicted activity.
The scores assigned to the explanations of LLMe2e are just slightly lower than the ones assigned to LLMExplainer. Overall, these results suggest that when no labeled data is available, LLMe2e is a good option since it provides reasonable recognition rates and it offers positively valued explanations. On the other hand, when training data is available, a combined technique with LLMExplainer applied to the output of a data-driven XAR method can lead to higher recognition rates and even more appreciated explanations. However, in the next section, we will discuss that, even though LLMExplainer explanations were highly appreciated, they may significantly increase the risk of over-reliance.
5. Drawbacks and Risks of Using LLMs to Generate Explanations
5.1. Over-reliance
Even though XAR may enable non-expert users to better understand the model’s predictions and increase their trust, natural language explanations that speculate over the simple facts detected by sensor data may lead to the over-reliance problem. Indeed, a user may excessively trust the system because it generates convincing explanations.
While this problem also occurs with existing approaches like DeXAR, it may be worsened when using LLMs, which are known to produce text that seems plausible but not always necessarily accurate (Hicks et al., 2024).
Considering the methods proposed in this paper, LLMe2e has a complete knowledge of the input data and of the ADL classification process to generate explanations, potentially offering a more accurate explanation of the decision process. On the other hand, LLMExplainer only has to generate an explanation given the most important features obtained from a data-driven classifier. We observed that LLMExplainer, in some instances, tends to generate explanations that are linguistically good but not supported by the observed sensor data. Hence, even though LLMExplainer produces appealing explanations highly rated in our surveys, these can be sometimes misleading, fostering an illusion of certainty and reliability. This problem becomes relevant when considering wrong predictions, as illustrated in the examples of Table 5.
Table 5. Examples of misclassified activity and over-reliance risks
Ground Truth DeXAR LLMExplainer on DeXAR LLMe2e
Watching TV I predicted the activity TAKING MEDICINES because the subject has repeatedly performed hand gestures and he has been in the dining room. I predicted the activity of TAKING MEDICINES because the subject showed dynamic hand movements for a short period, which suggests they were likely handling something, such as a medicine bottle or a pill. Additionally, the presence in the dining room during the same time indicates that this activity took place in a common area where people often take their medications. The combination of these factors strongly supports the prediction that the subject was taking their medicines. I predicted the activity PREPARING COLD MEAL because the subject moved from the kitchen to the dining room while showing dynamic hand movements, suggesting they were preparing something to eat
Eating I predicted the activity CLEARING TABLE because the subject has repeatedly performed hand gestures while being sitted in the dining room chair, and he has been in the dining room. I predicted the activity of CLEARING THE TABLE because the subject showed dynamic hand movements, indicating they were actively using their hands during the task. This activity was accompanied by pressure on the dining room chair, suggesting that the subject was seated while performing the task. Additionally, the presence in the dining room during this time further supports that the subject was engaged in clearing the table. Overall, the combination of these factors points to the activity of clearing the table. I predicted the activity EATING mainly because the subject was seated in the dining room and actively using their hands, which is consistent with eating.
Preparing hot meal I predicted the activity TAKING MEDICINES because the subject was sitting on the living room couch while television was on, and he has been in the living room for less than a minute. I predicted the activity of TAKING MEDICINES because the subject was present in the living room during this time and the living room chair was being used. Additionally, the TV was on, which often indicates a relaxed environment where someone might take their medicines. The combination of these factors suggests that the subject was likely engaged in this activity. I predicted the activity WATCHING TV mainly because the subject was sitting in the living room with the TV on and was making hand movements, which suggests they were interacting with the TV or a remote control.
While DeXAR and LLMExplainer use the same input to generate explanations, LLMExplainer generates significantly longer explanations that aim at convincing the reader that the relevant sensor states obtained by the XAR method are correlated with the predicted activity. Despite the system prompts of LLMe2e and LLMExplainer contain the same instructions for generating explanations, LLMe2e is also aware of the whole classification process. In our experiments this always led to more compact explanations that look less convincing on misclassifications, thus mitigating the over-reliance problem.
Even though it is not possible to completely eliminate this issue, some mitigation strategies can be applied. The first consists of integrating instructions in the system prompt aimed at avoiding motivations that stick as closely as possible to what is reported in the user message, thus not inventing nonexistent correlations. This strategy is already partially present in our methodology, in fact, as specified in Section 3.3.1 the system prompt includes the wording ”do not add inferences that are not directly supported by data”. Nevertheless, we believe that this strategy deserves further investigation and may lead to further mitigation of the problem we are observing.
Another possible and complementary strategy consists of passing to LLMExplainer also the input of the XAR model (i.e. the window of sensor events). In this way the LLM can use this additional information to produce a more factual explanation. Such approach will also produce an LLM-based prediction that could be combined with the one from the XAR classifier.
An additional mitigation strategy could be isolating and removing windows corresponding to transitions between activities. Indeed, as usually happens in HAR systems, a significant portion of the misclassifications occur while the subject transitions from one activity to another. During these transitions, the sensor events do not reflect any of the main target activities. Since the datasets we considered in this work do not have a label for transitions (e.g., “Other” or “Idle classes) the classifier is always forced to choose among one of the target activities, many of the misclassifications occur for this reason. As a consequence, the most important semantic states described in the explanations of these misclassifications are inconsistent with the predicted activity. While this phenomenon occurs for any XAR system, its impact may be even more negative for LLM-based approaches, where the model aims at creating convincing explanations for semantic states that are, however, inconsistent with the predicted activity. Hence, being able to isolate transitions could, in addition to improving classification performance, further help in reducing the number of misleading explanations.
Finally, a further strategy is using an agent-based approach. In this case, an LLM agent would be tasked with verifying that each rationale in the explanation is supported by the data and possibly providing an estimate of the risk of over-reliance.
5.2. Hallucinations in LLMs
Hallucinations in LLMs refer to instances where the model generates plausible but incorrect or fabricated information (Huang et al., 2023). In our domain, hallucinations can occur in activity classification and explanation generation.
Considering classification, hallucinations may lead to an incorrect activity, similar to mispredictions of traditional data-driven models. Additionally, hallucinations can manifest as outputs that do not comply with predefined instructions, such as predicting an activity outside the given set or producing results in an incorrect format. However, this second scenario never occurred in our experiments.
Considering explanations, hallucinations may introduce speculative or inconsistent reasoning. For instance, an LLM might invent correlations between sensor data to justify the predicted activity, creating an explanation that is linguistically coherent but not factually accurate. These cases can exacerbate over-reliance issues discussed in Section 5.1. In theory, LLMs may also generate entirely fabricated explanations poorly linked to the input, but we did not observe such extreme cases in our experiments. Nonetheless, the potential risk of hallucinations remains a challenge.
5.3. Applicability
Our experiments show that LLMs encode knowledge about ADLs. However, there are scenarios in which LLMe2e would fail. Indeed, we believe that LLMe2e can lead to good results only when the sensing infrastructure can properly capture the interaction of the subject with the environment. Such sensing infrastructures are commonly considered for healthcare monitoring applications (Grammatikopoulou et al., 2024) relying on cost-effective off-the-shelf devices.
On the other hand, considering datasets like the well-known CASAS (Cook et al., 2012), most of the sensors used in the homes are PIR sensors only detecting the movements of the inhabitants with all the limits that these sensors are known to have. Even though supervised data-driven models can capture motion patterns driven by ground truth (with the negative consequence of struggling in generalizing across different environments), motion sensors alone can not precisely reveal the actual activity performed by the subject. For this reason, LLMs struggle in mapping windows of motion sensor data into activities since it is challenging to use common-sense knowledge to determine ADLs from this data. Figure 16 shows an example of the step-by-step “reasoning” (through CoT) done by LLMe2e on a window extracted from the CASAS Milan dataset.
Figure 16. The Chain of Thought reasoning process of LLMe2e on the CASAS Milan dataset. The ground truth activity is work.
We adopted the segmentation window suggested in (Arrotta et al., 2022a) (i.e., 360 seconds) and we used the house floor plan to determine the semantic states generated by the deployed sensors. While the ground truth of this window is work, the LLM predicted kitchen activity since it considered, among the many fired motion sensors, the ones triggered in the kitchen as more important.
Based on the above considerations, it might not be worth spending money and resources to use LLMs when considering smart home deployments mostly based on PIR sensors.
5.4. Cost and privacy issues
In our experimental evaluation, we used a third party cloud-based LLM that is known to perform very well on many tasks beyond ADL recognition. However, the adoption of these models in real deployments may have two drawbacks: costs and privacy issues.
Considering the costs, both LLMe2e and LLMExplainer require a remote call to the LLM APIs for each window. At the time of writing, a single window for LLMe2e is costing on average $0.0085\$$ when using gpt-4o. Considering windows of $16$ seconds with $0.8$ of overlap, as we did for the MARBLE dataset, operating the system continuously would cost approximately $230\$$ per day.
Open-source LLMs are a promising alternative to reduce these costs. However, the most accurate models still require powerful computing machines equipped with GPUs and substantial amounts of RAM, resulting in costs not only for the hardware, but also for energy consumption. Dedicated machines could be deployed in a telemedicine infrastructure to optimize the costs by performing ADL recognition for a wide number of subjects. It is unclear if, in the future, more lightweight language models (e.g., small models like Phi-4 (Abdin et al., 2024) or quantized versions of open-source models) will potentially run on smart home gateways with acceptable performance and accuracy. Open-source models could also mitigate privacy problems since data about subjects’ actions in the home and their activities could be processed by a trusted entity and not released to an untrusted third-party provider.
6. Conclusion and Future Work
This paper explored how to leverage LLMs for XAR, showing the possible benefits, but also evaluating the drawbacks and risks. Our LLMe2e method is particularly promising for real-world applications, where labeled data collection is not feasible. Indeed, its ability to recognize activities in a zero-shot setting (i.e. when no training data is available) combines with the generation of explanations considered as high quality according to our user surveys. The LLMExplainer method demonstrates how LLMs can potentially generate explanations for any XAR method, thus improving their understandability for non-expert users.
Our critical evaluation highlighted the main problems in generating explanations with LLMs and illustrated several mitigation strategies that may be taken. Overall, we believe that once these challenges are addressed, the benefits will greatly outweigh the drawbacks and LLMs will become a powerful tool for implementing XAR systems in the wild.
Among the many research directions that this work is opening, we plan to test open-source LLM-based solutions on a large unlabeled dataset that we are acquiring from smart homes as part of a collaboration project with neurologists [^4] https://ecare.unimi.it/pilots/serenade/.